# Nearest Neighbor CCP-Based Molecular Sequence Analysis

### v1

@@ Abstract
Molecular sequence analysis is crucial for comprehending several biological processes, including protein-protein interactions, functional annotation, and disease classification. The large number of sequences and the inherently complicated nature of protein structures make it challenging to analyze such data. Finding patterns and enhancing subsequent research requires the use of dimensionality reduction and feature selection approaches. Recently, a method called Correlated Clustering and Projection (CCP) has been proposed as an effective method for biological sequencing data. The CCP technique is still costly to compute even though it is effective for sequence visualization. Furthermore, its utility for classifying molecular sequences is still uncertain. To solve these two problems, we present a Nearest Neighbor Correlated Clustering and Projection (CCP-NN)-based technique for efficiently preprocessing molecular sequence data. To group related molecular sequences and produce representative supersequences, CCP makes use of sequence-to-sequence correlations. As opposed to conventional methods, CCP doesn’t rely on matrix diagonalization, therefore it can be applied to a range of machine-learning problems. We estimate the density map and compute the correlation using a nearest-neighbor search technique. We performed molecular sequence classification using CCP and CCP-NN representations to assess the efficacy of our proposed approach. Our findings show that CCP-NN considerably improves classification task accuracy as well as significantly outperforms CCP in terms of computational runtime.
Index Terms:
Nucleotides, CCP, Spike sequence, Dimensionality Reduction
@@ I Introduction
Molecular sequences are a crucial part of the dynamic changes in sequence composition that control biological processes. Researchers have analyzed these molecular sequences and made progress toward a better understanding of physiology, biological development, and disease [1]. To understand various biological mechanisms and disorders, it is essential to understand how proteins interact with one another and carry out certain functions [2]. Unfortunately, because of their vastness, complexity, and lack of distinct patterns, it is still difficult to analyze thousands of molecular sequences at once. The development of medicines and treatments for human diseases is hampered by these obstacles. Deciphering the functions that proteins serve in various physiological and pathological circumstances is the goal of the field of proteomics [3]. Large sets of molecular sequences may now be made feasible by recent technical advancements to be analyzed to find patterns that could ultimately result in the creation of brand-new medications and vaccines [4]. However, processing such massive amounts of data necessitates the use of cutting-edge computing tools and statistical techniques to extract pertinent information.
@@
Additionally, It is well known that the data with a highly dimensional feature space will become sparse, which makes it difficult for statistical analysis to identify statistical significance and key factors. To facilitate prediction, analysis, and visualization, it is, therefore, preferable to minimize the dimensionality of high-dimensional data. Due to these challenges, a variety of dimensionality reduction (DR) techniques has been developed that can accurately reflect the inherent correlations in the original data on a low-dimensional space. There are several linear and non-linear DR methods have been proposed such as principal component analysis (PCA), Linear discriminant analysis (LDA), Multidimensional Scaling (MDS) [5], LargeVis [6] are a few linear dimensionality reduction methods. Whereas kernel PCA [7], Sammon mapping [8] and spectral embedding [9] are non-linear dimensionality reduction techniques.
@@
Correlated Clustering and Projection (CCP), a non-linear dimensionality reduction approach, computes the pairwise correlation matrix of samples and imposes a cutoff distance to prevent the global summation during the projection to increase computational efficiency [10, 11]. CCP has several benefits to offer, such as handling the dimensionality reduction of high sample sizes (because it avoids matrix diagonalization and instead solves a matrix to lower the dimensionality), employing statistical metrics like covariances to quantify the high-level dependence between random feature vectors [12], and it can be used in conjunction with a frequency-domain method for secondary dimensionality reduction to improve the preservation of data’s global structures and increase accuracy [13].
@@
This research is focused on creating a pipeline for analyzing molecular sequences using a method based on Nearest neighbor CCP-NN to preprocess the data and produce condensed representations that accurately reflect the original sequences. Here instead of computing the distance between data points, we use Nearest Neighbor (NN) to compute the nearest neighbor distance. We propose a method we name CCP-NN which is based on CCP to give an extra edge over the original CCP. The main contributions of our research are as follows:
@@
    We demonstrate the effectiveness of CCP and CCP-NN by evaluating their performance in molecular sequence classification tasks. Our results indicate that CCP-NN significantly enhances the accuracy of classification compared to other methods.
@@
    We offer a thorough framework for examining molecular sequence data by utilizing the capabilities of CCP. Our method facilitates effective feature selection, dimensionality reduction, and visualization.
@@ II Related Work
Sequence classification is a well-researched issue in bioinformatics [14]. A phylogenetic approach is frequently used in more conventional techniques of analyzing sequencing data [15], however, they are not scalable due to higher computational cost. To counter the issue, some machine learning (ML) approaches, including alignment-based [16, 17] and alignment-free [18] embedding approaches have become popular for ML tasks such as classification and clustering. Due to the extremely high dimensionality of the feature vector, these techniques do, however, also have scalability issues. The classification of biological sequences also makes use of the kernel matrix technique [19]. The Wasserstein distance (WD) is used in [20] to extract the features. Some efforts have been made to improve computational performance, such as Locality Sensitive Hashing (LSH) [21], which can train models faster and more accurately. The hash function in [22] is used to generate an approximate word embedding for language processing. However, collisions might occur in the resulting vectors, which reduces the embedding’s effectiveness. The use of bloom filters for mistake correction in raw read data to aid a de novo assembly was demonstrated by authors in [23]. However, these methods are prone to cause loss of information. Some of the most popular methods are principal component analysis (PCA) [24], Multidimensional Scaling [5], and LargeVis [6]. The curse of dimensionality and difficulties with the analysis of outliers are another issue [25]. When it comes to data noise, missing data, and poor-quality data, it can be quite unstable [25]. Sequence analysis is still difficult to perform despite extensive effort because of the high dimensionality and quantity of data [26]. Therefore, methods that can manage the dimensionality reduction of high sample numbers and do not involve matrix diagonalization are required.
@@ III Proposed Approach
We divide this section into two parts where we first discuss the original Correlated Clustering and Projection (CCP) method proposed by [10, 11]. After that, we discuss the proposed method, called Nearest Neighbor CCP-NN in detail.
@@ III-A Correlated Clustering and Projection (CCP) Algorithm
The Correlated Clustering and Projection (CCP) algorithm [10, 11] is a data clustering and dimensionality reduction technique that identifies and groups correlated features within a high-dimensional dataset. The algorithm operates by partitioning the features into clusters based on their correlation patterns and then projecting the data onto the subspace spanned by the identified clusters. The purpose of this algorithm is to capture the underlying structure of the data by focusing on feature subsets that exhibit strong correlations, thereby facilitating meaningful analysis and visualization. Given a dataset with $N$ samples and $M$ features represented by the matrix $\mathbf{X}$, the CCP proceeds as follows:
@@
The next step is to select a subset of features for clustering based on the variance. The algorithm chooses the top $numCutoff$ features, which is a user-defined parameter representing the percentage of non-zero variance features to retain.
@@
The selected features are clustered using the K-Means algorithm with $n\_components-1$ clusters, where $n\_components$ (a hyperparameter) is the desired number of clusters.
@@
The features that were not assigned to any cluster due to low variance are grouped into a separate cluster. This makes a new cluster that contains the remaining features.
@@
In Algorithm 1, the CCP performs several steps to compute the correlation and estimate the density map based on the given inputs and parameters. It begins by calculating the pairwise distances between the selected features. If the transformation flag is set to true, it calculates the distances between the features in the input data and the reference data. Otherwise, it calculates the distances between the features in the reference data. Next, if the scaling factor is not already calculated for the specified component, it proceeds to compute the average minimum distance, which is important for scaling the density estimation. Similarly, if the cutoff value is not already set for the specified component, it computes the average and standard deviation (SD) of the pairwise distances. The cutoff is then defined as the average plus three times the SD. It helps determine the threshold beyond which correlation values are considered negligible. The algorithm calculates the scaling factor by multiplying the user-defined scaling parameter with the previously computed average minimum distance. The scaling factor is used to adjust the scale of the density estimation. Finally, the algorithm estimates the density map, also known as the correlation, based on the calculated scaling factor and cutoff value. This estimation is done by computing the density of the pairwise distances using a density estimation function. The resulting density map represents the correlation between the selected features.
1:  computeCorrelations($index\_component$, $index\_Feat$, $X$, $transform$)
2:  if $transform$ then
3:     corr $\leftarrow$ pairwise_distances(X[:, index_Feat], self.X[:, index_Feat], self.metric)
@@
[…]
@@
[…]
@@
[…]
@@
14:  end if
15:  $Scale\leftarrow\text{self.scale}\times\text{self.avgmindist}[idx\_comp]$
@@
17:  $corr\leftarrow\textsc{computeDensity}(corr,Scale,cuttOffVal)$
18:  Return $corr$
@@ III-B Nearest Neighbors Based CCP
In the nearest neighbor (NN) version of CCP (our proposed method), all steps from 1 to 7 are followed from the original CCP as described in the above subsection. The main modification is made in Step 5, where we compute the density map using the NN algorithm for efficient and fast computation of the density map. The pseudocode for computing the density map is given in Algorithm 2, where the NearestNeighborComputeCorr function incorporates the use of an NN search technique, specifically the AnnoyIndex data structure [27], to calculate the correlation and estimate the density map. The steps involved in this process are as follows: An AnnoyIndex is created with the specified metric, and the features from the reference data are added to the index. The function checks if the transformation flag is set to true then we find a correlation by vector else if it is set to false, we find a correlation by item.
1:  NearestNeighborComputeCorr$(idx\_component,idx\_Feat,X,transform)$
@@
6:  $idx.build(-1)$
7:  if $transform$ then
@@
[…]
@@
20:  end if
21:  $Scale\leftarrow\text{self.scale}\times\text{self.avgmindist}[idx\_component]$
@@
23:  $corr\leftarrow\textsc{computeDensity}(corr,Scale,cuttOffVal)$
24:  return $corr$
@@ III-C1 CCP
To begin, the computation of variance along each feature axis takes $\mathcal{O}(N\cdot M)$, where $N$ is the number of samples and $M$ is the number of features. Sorting the variance values to select the top $f\ (\text{where}\ f\leq M)$ features takes $\mathcal{O}(f\log f)$. Following this, the K-Means clustering step typically depends on the number of iterations $n_{\text{iter}}$, the number of features $f$, the number of clusters $n_{c}$, and the number of samples $N$. This step has a time complexity of $\mathcal{O}(n_{c}\cdot f\cdot n_{\text{iter}}\cdot N)$.
@@
Furthermore, in order to compute the density map to capture the correlation between features within a cluster, pairwise distance calculation is required, which is the most computationally expensive operation. If $f_{i}$ is the number of features selected for the $i$-th cluster, the pairwise distance calculation takes $\mathcal{O}(N^{2}\cdot f_{i})$. This is because pairwise distance calculation compares each of the $N$ samples with every other sample across $f_{i}$ features. Summing over all $n_{c}$ components, the total time complexity becomes:
@@
where $f=\sum_{i=1}^{n_{c}}f_{i}$ is the total number of features across all components. The space complexity associated with this procedure is $\mathcal{O}(N^{2})$. Therefore, the overall complexity is;
@@ III-C2 CCP-NN
On the other hand, if we were to use Approximate Nearest Neighbor (ANN) via Annoy Index as a proxy for pairwise distance computation, we would significantly benefit in terms of computation time. Building an Annoy Index takes $\mathcal{O}(N\log N\cdot f)$, because the algorithm builds a forest of random projection trees. Querying the Annoy Index for nearest neighbors is approximately $\mathcal{O}(\log N)$ per query, and since this is repeated for all $N$ samples, the complexity for querying is $\mathcal{O}(N\log N)$.
@@
Clearly, CCP-NN has an advantage over CCP in terms of speed and memory requirement.
@@ III-D Convergence Analysis
where $p(x_{i},\hat{N}_{k}(x_{i}))$ denotes the density estimate using the true nearest-neighbor. We take expectation on both sides and the first term of R.H.S in the inequality above can be bounded by $\mathcal{O}(\epsilon)$ due to the accuracy of the Annoy Index. The second term can be bounded using standard kernel density estimation convergence results [28].

### v2

@@ Abstract
Molecular sequence analysis is crucial for understanding several biological processes, including protein-protein interactions, functional annotation, and disease classification. The large number of sequences and the inherently complicated nature of protein structures make it challenging to analyze such data. Finding patterns and enhancing subsequent research requires the use of dimensionality reduction and feature selection approaches. Recently, a method called Correlated Clustering and Projection (CCP) has been proposed as an effective method for biological sequencing data. The CCP technique remains computationally expensive, despite its effectiveness for sequence visualization. Furthermore, its utility for classifying molecular sequences is still uncertain. To solve these two problems, we present a Nearest-Neighbor Correlated Clustering and Projection (CCP-NN)-based technique for efficiently preprocessing molecular sequence data. To group related molecular sequences and produce representative supersequences, CCP makes use of sequence-to-sequence correlations. As opposed to conventional methods, CCP does not rely on matrix diagonalization, therefore, it can be applied to a range of machine-learning problems. We estimate the density map and compute the correlation using a nearest-neighbor search technique. We perform a molecular sequence classification using CCP and CCP-NN representations to assess the efficacy of our proposed approach. Our findings show that CCP-NN considerably improves the accuracy of the classification task and significantly outperforms CCP in terms of computational runtime.
@@ I Introduction
Molecular sequences are a crucial part of the dynamic changes in sequence composition that control biological processes. Researchers have analyzed these molecular sequences and have made progress toward a better understanding of physiology, biological development, and disease [1]. To understand various biological mechanisms and disorders, it is essential to understand how proteins interact with each other and carry out certain functions [2]. Unfortunately, because of their vastness, complexity, and lack of distinct patterns, it is still difficult to analyze thousands of molecular sequences at once. These obstacles hamper the development of medicines and treatments for human diseases. Deciphering the functions that proteins serve in various physiological and pathological circumstances is the goal of the field of proteomics [3]. Large sets of molecular sequences can now be made feasible by recent technical advancements to be analyzed to find patterns that could ultimately create brand new medications and vaccines [4]. However, processing such massive amounts of data requires the use of cutting-edge computing tools and statistical techniques to extract pertinent information.
@@
Furthermore, it is well known that data with a highly dimensional feature space will become sparse, making it difficult for statistical analysis to identify statistical significance and key factors [5]. To facilitate prediction, analysis, and visualization, it is preferable to minimize the dimensionality of high-dimensional data. Due to these challenges, a variety of dimensionality reduction (DR) techniques have been developed that can accurately reflect the inherent correlations in the original data in a low-dimensional space. Several linear and non-linear DR methods have been proposed, such as principal component analysis (PCA), linear discriminant analysis (LDA), Multidimensional Scaling (MDS) [6], LargeVis [7] is a few linear dimensionality reduction methods. Whereas kernel PCA [8, 9, 10], Sammon mapping [11] and spectral embedding [12] are non-linear dimensionality reduction techniques.
@@
Correlated Clustering and Projection (CCP), a non-linear dimensionality reduction approach, computes the pairwise correlation matrix of samples and imposes a cutoff distance to prevent the global summation during the projection to increase computational efficiency [13, 14]. CCP has several benefits to offer, such as handling the dimensionality reduction of high sample sizes (because it avoids matrix diagonalization and instead solves a matrix to lower the dimensionality), employing statistical metrics such as covariances to quantify the high-level dependence between random feature vectors [15], and it can be used in conjunction with a frequency domain method for secondary dimensionality reduction to improve the preservation of data’s global structures and increase accuracy [16].
@@
This research is focused on creating a pipeline for analyzing molecular sequences using a method based on Nearest Neighbor CCP-NN to preprocess the data and produce condensed representations that accurately reflect the original sequences. Here, instead of computing the distance between data points, we use Nearest Neighbor (NN) to compute the nearest neighbor distance. We propose a method, which we name CCP-NN, that is based on CCP to give an extra edge over the original CCP. The main contributions of our research are as follows:
@@
    We demonstrate the effectiveness of CCP and CCP-NN by evaluating their performance in molecular sequence classification tasks. Our results indicate that CCP-NN significantly improves the accuracy of classification compared to other methods.
@@
    We offer a thorough framework for examining molecular sequence data using the capabilities of CCP. Our method facilitates effective feature selection, dimensionality reduction, and visualization.
@@ II Related Work
Sequence classification is a well-researched issue in bioinformatics [17, 18, 19]. A phylogenetic approach is frequently used in more conventional techniques of analyzing sequencing data [20], however, they are not scalable due to higher computational cost. To counter the issue, some machine learning (ML) approaches, including alignment-based [21, 22, 23, 24] and alignment-free [25] embedding approaches, have become popular for ML tasks such as classification and clustering. Due to the extremely high dimensionality of the feature vector, these techniques do, however, also have scalability issues. The classification of biological sequences also makes use of the kernel matrix technique [26]. The Wasserstein distance (WD) is used in [27] to extract the features. Some efforts have been made to improve computational performance, such as Locality Sensitive Hashing (LSH) [28], which can train models faster and more accurately. The hash function in [29] is used to generate an approximate word embedding for language processing. However, collisions might occur in the resulting vectors, which reduces the embedding’s effectiveness. The use of bloom filters for the correction of errors in raw read data to aid a de novo assembly was demonstrated by the authors in [30]. However, these methods are prone to information loss. Some of the most popular methods are principal component analysis (PCA) [31], Multidimensional Scaling [6], kernel selection methods [32], and LargeVis [7]. The curse of dimensionality and the difficulties with the analysis of outliers are another issue [33]. When it comes to data noise, missing data, and poor-quality data, it can be quite unstable [33]. Sequence analysis is still difficult to perform despite extensive effort due to the high dimensionality and quantity of data [34]. Therefore, methods that can manage the dimensionality reduction of high sample numbers and do not involve matrix diagonalization are required.
@@ III Proposed Approach
We divide this section into two parts, where we first discuss the original Correlated Clustering and Projection (CCP) method proposed by [13, 14]. After that, we discuss the proposed method, called Nearest Neighbor CCP-NN, in detail.
@@ III-A Correlated Clustering and Projection (CCP) Algorithm
The Correlated Clustering and Projection (CCP) algorithm [13, 14] is a data clustering and dimensionality reduction technique that identifies and groups correlated features within a high-dimensional dataset. The algorithm operates by partitioning the features into clusters based on their correlation patterns and then projecting the data onto the subspace spanned by the identified clusters. The purpose of this algorithm is to capture the underlying structure of the data by focusing on feature subsets that exhibit strong correlations, thus facilitating meaningful analysis and visualization. Given a dataset with $N$ samples and $M$ features represented by the matrix $\mathbf{X}$, the CCP proceeds as follows:
@@
The next step is to select a subset of features for clustering based on variance. The algorithm chooses the top $numCutoff$ features, which is a user-defined parameter representing the percentage of non-zero variance features to retain.
@@
The selected features are clustered using the K-Means algorithm with n_components1n\_components-1 clusters, where n_componentsn\_components (a hyperparameter) is the desired number of clusters.
@@
The features that were not assigned to any cluster due to low variance are grouped into a separate cluster. This creates a new cluster that contains the remaining features.
@@
In Algorithm 1, the CCP performs several steps to compute the correlation and estimate the density map based on the inputs and parameters given. It begins by calculating the pairwise distances between the selected features. If the transformation flag is set to true, it calculates the distances between the features in the input data and the reference data. Otherwise, it calculates the distances between the features in the reference data. Next, if the scaling factor is not already calculated for the specified component, it proceeds to compute the average minimum distance, which is important for scaling the density estimation. Similarly, if the cutoff value is not already set for the specified component, it computes the average and standard deviation (SD) of the pairwise distances. The cutoff is then defined as the average plus three times the SD. It helps determine the threshold beyond which correlation values are considered negligible. The algorithm calculates the scaling factor by multiplying the user-defined scaling parameter with the previously computed average minimum distance. The scaling factor is used to adjust the scale of the density estimation. Finally, the algorithm estimates the density map, also known as the correlation, based on the calculated scaling factor and cutoff value. This estimation is done by computing the density of the pairwise distances using a density estimation function. The resulting density map represents the correlation between the selected features.
1: computeCorrelations(index_component,index_Feat,X,transform)(index\_component,index\_Feat,X,transform)
2:if transformtransform then
3:  $corr\leftarrow\textsc{pairwise\_distances}(X[:,index\_Feat],\text{self.X}[:,index\_Feat],\text{self.metric})$
@@
5:  $corr\leftarrow\textsc{pairwise\_distances}(\text{self.X}[:,index\_Feat],\text{self.metric})$
@@
7:if $\text{self.avgmindist}[index\_component]==0$ then
8:  $\text{self.avgmindist}[index\_component]\leftarrow\textsc{computeAvgMinDistance}(corr)$
@@
10:if $\text{self.cutoff}[index\_component]==0$ then
@@
13:  $\text{self.cutoff}[index\_component]\leftarrow avg+3\times std$
14:end if
@@
[…]
18:return corrcorr
@@ III-B Nearest Neighbors Based CCP
In the nearest neighbor (NN) version of CCP (our proposed method), all steps from 1 to 7 are followed from the original CCP as described in the above subsection. The main modification is made in Step 5, where we compute the density map using the NN algorithm for efficient and fast computation of the density map. The pseudocode for computing the density map is given in Algorithm 2, where the NearestNeighborComputeCorr function incorporates the use of an NN search technique, specifically the AnnoyIndex data structure [35], to calculate the correlation and estimate the density map. The steps involved in this process are as follows: An AnnoyIndex is created with the specified metric, and the features from the reference data are added to the index. The function checks if the transformation flag is set to true, then we find a correlation by vector; else if it is set to false, we find a correlation by item.
1:NearestNeighborComputeCorr(idx_component,idx_Feat,X,transform)(idx\_component,idx\_Feat,X,transform)
@@
6:idx.build(1)idx.build(-1)
7:if transformtransform then
@@
13:if $\text{self.avgmindist}[idx\_component]==0$ then
14:  $\text{self.avgmindist}[idx\_component]\leftarrow\textsc{computeAvgMinDistance}(corr)$
@@
19:  $\text{self.cutoff}[idx\_component]\leftarrow avg+3\times std$
20:end if
@@
[…]
24:return corrcorr
@@ III-C Algorithm Complexity
In this section, we discuss the time and space complexity of CCP and CCP-NN.
@@ III-C1 CCP
[…]
@@
Furthermore, in order to compute the density map to capture the correlation between features within a cluster, a pairwise distance calculation is required, which is the most computationally expensive operation. If $f_{i}$ is the number of features selected for the $i$-th cluster, the pairwise distance calculation takes $\mathcal{O}(N^{2}\cdot f_{i})$. This is because the pairwise distance calculation compares each of the NN samples with every other sample across the $f_{i}$ features. Summing over all $n_{c}$ components, the total time complexity becomes:
@@
where $f=\sum_{i=1}^{n_{c}}f_{i}$ is the total number of features in all components. The space complexity associated with this procedure is $\mathcal{O}(N^{2})$. Therefore, the overall complexity is;
@@ III-C2 CCP-NN
On the other hand, if we were to use Approximate Nearest Neighbor (ANN) via the Annoy Index as a proxy for pairwise distance computation, we would benefit significantly in terms of computation time. Building an Annoy Index takes $\mathcal{O}(N\log N\cdot f)$, because the algorithm builds a forest of random projection trees. Querying the Annoy Index for nearest neighbors is approximately $\mathcal{O}(\log N)$ per query, and since this is repeated for all $N$ samples, the complexity of querying is $\mathcal{O}(N\log N)$.
@@
Clearly, CCP-NN has an advantage over CCP in terms of speed and memory requirements.
@@ III-D Convergence Analysis
where $p(x_{i},\hat{N}_{k}(x_{i}))$ denotes the density estimate using the true nearest-neighbor. We take the expectation on both sides, and the first term of the R.H.S in the inequality above can be bounded by $\mathcal{O}(\epsilon)$ due to the accuracy of the Annoy Index. The second term can be bounded using the convergence results of the standard kernel density estimation [36].
@@
Our method produces a number of features equal to 20, which are selected using standard cross-validation. Each CCP component corresponds to one output feature, where each feature represents the aggregated correlation/density within a specific subset of the original feature space. The original high-dimensional input is therefore reduced to 20 CCP-derived features that capture the local density structure of different feature groupings.